A pharmaceutical company Eureka Bio has discovered a Corona vaccine that can be produced at constant marginal cost of R10. The company has entered into offtake dosage agreements with Country A and B. Country A has a dosage demand of QA = 200 - PA, and Country B has dosage demand QB = 160 - PB.


(c) If World Health Organization introduces a regulation on the price of dosages, calculate the price, profits and dosages that Eureka can charge. [5]  


1
Expert's answer
2021-07-11T22:10:03-0400

"MC=P\\P=10\\QA=200-10=190\\QB=160-(-10)170\\Q=QA+QB=360"